OpenStudy (loser66):

How to solve: \(x = p(1-e^{-kx})^2\) where k and p are constant Please, help

OpenStudy (loser66):

@dan815

OpenStudy (retireed):

What are you solving for?

OpenStudy (trojanpoem):

He said k , p are constants so no other variable than x

OpenStudy (loser66):

for x , I believe

OpenStudy (retireed):

Thank's looks like a lot of algebra Need to get some paper.

OpenStudy (michele_laino):

please, try to use the Taylor expansion of \(\Large e^{-kx}\): \[\huge {e^{ - kx}} \approx 1 - kx + \frac{{{k^2}{x^2}}}{2}\]

OpenStudy (loser66):

If so, we use power series for \(e^{-kx}= \sum_{n=0}^\infty \dfrac{(-kx)^n}{n!}\) . Is it a mess?

OpenStudy (retireed):

That is an excellent idea, because simple algebra won't solve this equation.

OpenStudy (michele_laino):

Ithink it is a way in order to get an algebraic equation, which can be easily solved

OpenStudy (michele_laino):

oops.. I think...

OpenStudy (loser66):

\(1- e^{-kx} )^2 = 1 - 2e^{-kx}+e^{-2kx}= e^{-kx}(e^{kx}-2+e^{-kx})\) and we know that \(2cos(kx) = e^{kx}+e^{-kx}\) hence it becomes \(2e^{-kx}(cos(kx) -1)\)

OpenStudy (kainui):

A quick guess is x=0

OpenStudy (michele_laino):

That's right!! @Kainui

OpenStudy (kainui):

You can confirm that it's the only solution by taking the derivative and seeing that it's going to be the only solution.

OpenStudy (kainui):

I mean because the slopes of the functions will have only one intersection and all that... I'm a little tired haha hard to explain myself.

OpenStudy (michele_laino):

furthermore, since we are searching for an intersection between a straight line \(y=x\) and the exponential function, we can consider only the linear term of the Taylor expansion, so we get the subsequent equation: \[\huge x = p\left( {{k^2}{x^2}} \right)\]

OpenStudy (loser66):

I got what you mean. Thanks a lot.

OpenStudy (michele_laino):

therefore, we get the subsequent solutions: \[\huge x = 0,\quad x = \frac{1}{{p{k^2}}}\]

OpenStudy (kainui):

Here's a trick to remove a constant: \(y=-kx\) \[y=-kp(1-e^y)^2\] \[-kp=c\] is the new arbitrary constant: \[y=c(1-e^y)^2\]

OpenStudy (loser66):

and then?